# Finding 3D Positions of Distant Objects from Noisy Camera Movement and Semantic Segmentation Sequences

### v1

@@ Abstract
3D object localisation based on a sequence of camera measurements is essential for safety-critical surveillance tasks, such as drone-based wildfire monitoring. Localisation of objects detected with a camera can typically be solved with dense depth estimation or 3D scene reconstruction. However, in the context of distant objects or tasks limited by the amount of available computational resources, neither solution is feasible. In this paper, we show that the task can be solved using particle filters for both single and multiple target scenarios. The method was studied using a 3D simulation and a drone-based image segmentation sequence with global navigation satellite system (GNSS)-based camera pose estimates. The results showed that a particle filter can be used to solve practical localisation tasks based on camera poses and image segments in these situations where other solutions fail. The particle filter is independent of the detection method, making it flexible for new tasks. The study also demonstrates that drone-based wildfire monitoring can be conducted using the proposed method paired with a pre-existing image segmentation model.
@@ I Introduction
This work addresses the problem of locating distant objects from a series of camera-based detections from known locations and orientations. At a glance, the problem of locating target objects based on multiperspective imagery seems fairly well-addressed. The earlier proposed methods typically operate as optimisation problems where the object locations of individual keypoints are determined using a set of images and known camera poses. Alternatively, full scene 3D reconstruction or dense depth estimation methods have been used for similar problems.
However, the task at hand, presented in this work, is specific to far-away objects which have been detected with separate image segmentation models, such as neural networks which inherently produce noisy segments. Other sources of noise occur from possible target object dynamics and the camera-pose estimation methods. In addition, due to the nature of the target detection models, direct correspondence between features from consecutive frames does not hold.
The motivation for this work originates from drone-based wildfire detection, in which the position of the drone-carried camera is estimated using GNSS measurements and known dynamics of the drone camera setup. Our previous work showed that wildfire smoke can be detected from almost ten kilometres away using only drone-carried resources [Pesonen_2025_WACV]. Pairing the segmentation model with a lightweight target localisation method enables fully on-board wildfire detection and localisation. This enables the wildfire detection system to be used in areas of poor telecommunication where cloud-based computing can not be relied on. The sketch in Figure 1 illustrates a use case of a UAV scanning for wildfires (sketch by DALL-E 3) with masked RGB images (real data).
@@
To combat the noise induced by the camera and image observation systems, we focus on Bayesian filters. They enable reliable modelling of various noise sources, and with particle filters in particular, more complex observation and target dynamics can be modelled. The use of particle filters in camera-based 3D localisation literature has been limited. Specifically, for locating distant objects from a moving camera, there appears to be a lack of extensive research.
This work extends the literature on filter-based target object localisation from a moving camera by addressing the problem of distant objects detected by separate models. We propose using particle filters to iteratively improve the localisation and uncertainty estimation of the target object’s 3D position. Using both simulations and a drone-captured image sequence with GNSS-estimated camera positions, we show that the method can locate multiple target objects from moving camera segmentation sequences despite the various sources of noise.
II Related Work
The task at hand can be thought of as a form of camera-based multiview 3D reconstruction. The earlier developments of 3D object reconstruction from multiview imagery have been well documented by Hartley and Zisserman [Hartley2004]. These methods relied on point correspondence, well-estimated camera parameters, and bundle adjustments. Later developments, such as COLMAP [schoenberger2016sfm] and 3D Gaussian splatting [kerbl3Dgaussians], have greatly improved the structure from motion map generation and 3D modelling, respectively. Unfortunately, the methods still require finding a large number of point correspondences between frames, making them computationally heavy for iterative real-time localisation using edge devices, which is desirable in, for example, aerial robotic applications.
The 3D reconstruction methods have been poorly studied in the context of smaller and more distant objects. In such situations, any noise in the camera pose estimation causes larger errors in the final reconstruction. This suggests that Bayesian filtering-based solutions could offer feasible solutions. Besides the camera pose estimation errors, typical modern camera-based sensing solutions leverage neural network models for detecting key features, which complicates keypoint matching between frames.
@@ II Related Work
Even though it’s difficult to find direct correspondence for the problem in camera-focused literature, object localisation and tracking in 3D space is a largely studied problem in the context of other sensor modalities, particularly radars. Different types of multitarget tracking algorithms are well covered by a survey by Vo et al. [vo2015multitarget] and a later survey focusing on particle filter-based methods by Wang et al. [wang2017survey]. While new algorithms have been proposed since these surveys, they cover the basics in high detail and provide a good overview of the different approaches.
The use of particle filters for tracking multiple targets has been proposed at least as early as 2002 by Hue et al. [Hue2002TrackingMO], and for cameras in the 2D camera plane, a particle filter has been suggested for the problem as early as 2005 by Yang et al. [yang2005fast]. Ever since, the methods have evolved, particularly in the way that multiple objects are represented. A survey by Rao et al has covered techniques used in particle filters for tracking of visual targets [rao2013visual]. Examples of multitarget presentation methods include multiple filter methods, which were originally introduced for the problem of high dimensionality [djuric2007multiple]. The multiple filter methods split the task into multiple single-target filtering tasks. Using multiple filters introduces the problem of assigning the correct observations to each filter. This has been solved, for example, with the symbiotic particle filter [bugallo2010target] or by estimating unknown parameters using fused information from multiple particle filters [zhao2025fusion]. Another solution for multiple targets is presented by finite-set statistics developed by Ron Mahler [9100524], which are used for the PHD (probability hypothesis density) particle filters. These have also been applied to 2D multitarget tracking in camera images [maggio2007particle].
Tracking multiple objects using multiple sensors has also been the focus of many of the more recent works, also using random finite set presentations, as covered in a survey by Da et al. [da2021recent]. In the multisensor context, the use of cameras has been slightly more common as well. However, in the multicamera scenario, the problem is inherently different from that of a single camera, due to the possibility of using the discrepancies between the multiple sensor locations. Besides, the metrics in the literature have been focused on the 2D labels as with individual cameras [AMOSA2023126558].
@@ III-A Problem Setup
where $\mathbf{y}$ is the projected point in homogeneous coordinates, $\mathbf{K}$ is the intrinsic camera matrix, $\mathbf{M}$ is the extrinsic camera matrix, and $\mathbf{x}$ is the original 3D point in homogeneous coordinates. The homogeneous coordinates use an extra dimension to simplify the matrix operations, such that $y=(y_{1},y_{2},y_{3})^{\top}$ and $x=(x_{1},x_{2},x_{3},x_{4})^{\top}$, where the pixel coordinates of the projection are:
@@
The false positives in the simulated segments were generated by defining random rectangular sections of the image and setting these pixels equal to the real positive segments. These false-positive rectangles were defined by their height and width in pixel dimensions. They were generated randomly for each image, based on a false positive rate $\rho_{FP}$ and the maximum number of false positives $Max_{FP}$. Correspondingly, the false positives were removed from the following frames based on a false positive dismissal rate $\delta\rho_{FP}$. Unless the false positive was removed, it was kept for the subsequent images.
@@
III-B Single Target Particle Filter
We defined the filter such that the particles were initialised based on a traditional least squares estimate between the means of two observations, after a threshold number, $\tau_{min-obs}$, of consecutive positive observations.
@@ III-B Single Target Particle Filter
was done by defining the particles as Monte-Carlo samples of a three-dimensional Gaussian distribution, scaled in each dimension based on the distance between the camera and the observed target. Thus, the initial distribution took the form
$\boldsymbol{\pi}_{init}\sim\mathcal{N}(\mathbf{m}_{init},\mathbf{P}_{init}),$ (7)
where $\mathbf{m}_{init}$ is the middle point between the two camera rays, $r_{1}$ and $r_{2}$, computed using the inverse operation of the camera matrices, and the middle point of each observation in the camera frame. $\mathbf{P}_{init}$ is the covariance of the initial distribution, which was set at a constant that was determined by adjusting the model such that the initial distribution would be very likely to contain particles located at the target position while simultaneously remaining dense enough to not cause issues in convergence.
@@
was first taken after the initialisation. In it, the particle distribution was updated based on new observations. At each update step, the particles were injected with independent Gaussian noise, based on the distance between the point and the camera and an optimised constant, which meant that after each observation step, the full distribution could be expressed as a combination of N three-dimensional Gaussians, where N is the number of particles.
@@
required that positive pixels were observed. Given this, the pixel distances between the projected particles and the positive observations were compared. Each particle, projected inside the camera frame, was assigned a weight, $\mathbf{\omega}_{p}$, relative to the distance from the positive pixels such that:
$\mathbf{\omega}_{p}=e^{-\min((\mathbf{obs}-p_{proj})^{2})},$ (8)
where $\mathbf{obs}$ is an array of the positive observations and $p_{proj}$ is the projected particle in the pixel coordinates.
@@
took place after weighing each particle. The resampling was performed by using the weights as the probabilities of drawing the corresponding points from the set of particles. The same number of particles as the previous set was drawn in this manner, resulting in a new set where the highest weighted particles were repeated multiple times. This resulting distribution was then used in the next prediction step, where particles were first updated randomly, effectively doing Gaussian kernel smoothing for the distribution, before projecting them again to the camera frame. For the study, we used the bootstrap version of the particle filter, due to better initial results, meaning that the resampling was done at every observation.
III-C Multitarget Extension
Extending the particle filter for multiple targets is possible using various formulations as introduced in Section II. In this work, we solved the multiple target scenario using multiple particle filters. We opted for this particular method to avoid using an unnecessarily large distribution for the particle filter prediction steps, and instead, used a larger number of more focused distributions of separate particle filters.
In practice, the multiple target localisation was realised by computing the pixels at each observation that were located beyond a dynamic threshold, $\theta_{p-o}$, from the nearest projected particles. Given that these out-of-distribution (OOD) observations appeared at a certain number of consecutive observation steps, $\tau_{min-obs}$, a new particle filter was initialised based on the OOD pixels. The separate particle filters were also dismissed in case no positive observations were found for a similar threshold number of observation steps, $n_{\theta-dm}$.
@@ III-D Metrics
The performance of the proposed method was quantified using two measures, the root mean square error (RMSE) and negative log predictive density (NLPD), of which the RMSE was computed between the predicted particles, $\mathbf{p}_{i}$, and the mean of the target location, $\mathbf{m}_{t}$:
$\mathrm{RMSE}=\sqrt{\dfrac{\sum_{i=1}^{n_{p}}(\mathbf{p}_{i}-\mathbf{m}_{t})^{2}}{N}},$ (9)
where $n_{p}$ is the number of particles. This is equal to the Euclidean distance between the predicted and ground truth means. The NLPD was computed based on the likelihood that the target mean would be drawn from the distribution of particles as if they presented a normal distribution, such that:
$\mathrm{NLPD}=\dfrac{-\sum_{i=1}^{n_{p}}\log\ p(\mathbf{m}_{t}\mid\mathbf{p})}{n_{p}}.$ (10)
The RMSE presents an intuitive linear measure for the distance between the predicted target centre point and the true target centre point, while the NLPD better takes into account the uncertainty presented by the variation of the particles.
For both metrics, we report the minimum obtained value during the test, corresponding to the converged filter estimate, and for the RMSE, we also report the mean between 200 and 1000 metres of camera translation. The first 200 metres were dismissed due to being mostly dependent on the initialisation, and 1000 metres was the maximum translation in the experiments. For all results, the metrics were computed over an average of ten simulations to reduce noise caused by the randomness of the experiments.
@@ III-E Empirical Data
To validate that the performance of the proposed method holds for real-world applications, we tested the system with a drone-captured video sequence of a telecommunication mast. The position of the drone was recorded using differential GNSS, and the geolocation of the target object was known.
To capture the sequence, we used a DJI Matrice 350, equipped with an AR0234 camera and an Applanix APX-15 UAV GNSS. An NVIDIA Jetson Orin NX was used to record the data. We performed no calibration for the camera or the GNSS. This means that for the camera, only the parameters provided by the camera manufacturer were used [arducamAR0234Arducam]. The camera was used to record at full HD (1080 by 1920 pixels) resolution with a horizontal FOV of 90 and vertical FOV of 50.625. This means that the intrinsic parameters were set as $f_{x}=1200$, $f_{y}=1200$, $c_{x}=960$, and $c_{y}=540$. The lens was assumed to cause no distortions. The GNSS antenna was mounted approximately 30 centimetres away from the camera with an external IMU attached to the camera. Neither the boresight nor the lever arm of the mounting was taken into account for the experiments. This means that the camera pose estimation had at least a systematic error in the range of tens of centimetres in addition to any sensor-caused errors. The manufacturer reported RMSE for the position is 0.02 to 0.05 metres, for the roll and pitch 0.025 degrees, and for the heading 0.080 degrees [APX15UAV].
@@
$N_{T}$ Max $\nu_{rot}$ ($\circ$) Max $\nu_{t}$ (m) $\rho_{FP}$ $\delta\rho_{FP}$ Max FPFP $\rho_{FN}$ $\rho_{PFN}$ $\delta\rho_{PFN}$ RMSE min (m) RMSE 200-1k (m) NLPD min
1 0 0 0 - 0 0 0 - 37.81 140.57 14.42
1 0.1 0.5 0 - 0 0 0 - 36.93 141.04 14.31
1 0.1 0.5 0.1 0.2 3 0 0 - 47.54 168.19 14.44
1 0.1 0.5 0.1 0.2 3 0.1 0 - 52.44 168.19 14.44
1 0.1 0.5 0.1 0.2 3 0.1 0.1 0.2 80.00 198.03 18.10
3 0 0 0 0 0 0 0 - 171.56 264.87 15.80
3 0.1 0.5 0 0 0 0 0 - 158.25 239.19 16.79
3 0.1 0.5 0.1 0.2 3 0 0 - 213.41 296.38 16.99
3 0.1 0.5 0.1 0.2 3 0.1 0 - 231.05 361.40 46.17
3 0.1 0.5 0.1 0.2 3 0.1 0.1 0.2 265.05 484.00 25.74
The images were post-processed using a horizontal Sobel operation with image erosion and dilation. This way, we obtained a sequence of binary segments with a known target location and estimated camera positions, corresponding to a drone-based wildfire localisation task. The known target geolocation also enabled measuring the quality of the localisation in ground coordinates.

### v2

# Julius Pesonen    Arno Solin    Eija Honkavaara

@@ Abstract
3D object localisation based on a sequence of camera measurements is essential for safety-critical surveillance tasks, such as drone-based wildfire monitoring. Localisation of objects detected with a camera can typically be solved with specialised sensor configurations or 3D scene reconstruction. However, in the context of distant objects or tasks limited by the amount of available computational resources, neither solution is feasible. In this paper, we show that the task can be solved with either multi-view triangulation or particle filters, with the latter also providing shape and uncertainty estimates. We studied the solutions using 3D simulation and drone-based image segmentation sequences with global navigation satellite system (GNSS) based camera pose estimates. The results suggest that combining the proposed methods with pre-existing image segmentation models and drone-carried computational resources yields a reliable system for drone-based wildfire monitoring. The proposed solutions are independent of the detection method, also enabling quick adaptation to similar tasks.
@@ I Introduction
This work explores solutions for locating very distant objects from a series of camera-based detections from known locations and orientations. At a glance, the problem of locating target objects based on multiperspective imagery seems well-addressed. However, the task poses unique problems when applied to objects that might be located multiple kilometres away. Typical solutions to such tasks include specific sensor configurations relying on stereo cameras or time-of-flight sensors, such as lidar. Unfortunately, at such scales, stereo cameras would require huge baselines, and time-of-flight sensors would become unreliable due to their 3D spatial resolution worsening cubically by distance. Alternatively, full camera-based 3D reconstruction methods have been applied to such problems, but creating a 3D model of such a vast region becomes computationally inefficient when the goal is to determine the position of a single target or those of only a few target objects.
The motivation for this work originates from drone-based wildfire detection, in which the position of the drone-carried camera is estimated using GNSS measurements and known dynamics of the drone camera setup. Our earlier work [Pesonen_2025_WACV] showed that wildfire smoke can be detected from almost ten kilometres away using only drone-carried resources. Pairing the segmentation model with a lightweight target localisation method would enable fully on-board wildfire detection and localisation. This means that the wildfire detection system could be deployed in areas of poor telecommunication where cloud-based computing can not be relied on. The sketch in Figure 1 illustrates a use case of a UAV scanning for wildfires (sketch by DALL-E 3) with masked RGB images (real data).
@@
Besides the perception distance, smoke clouds as the perceived targets pose unique challenges due to their practically unlimited variety of shapes. To get a reliable estimate of the possible wildfire positions in ground coordinates, it would be beneficial to get a prediction of the shape of the smoke cloud and any possible uncertainties presented by the perception framework. A variety of sources for uncertainties are presented in the task, as even small errors in the camera pose estimation cause major differences in perceived positions as the sensing distances grow larger. With far enough objects, it also becomes practically impossible to see a large distant object from such a variety of angles that it could be perfectly enclosed in a single position. Thus, it’s beneficial to use methods which could present the full region of possible locations for the target object.
To study the possible solutions, we created a simple simulation where target objects are simplified as cubes. We also evaluated the presented methods using two sequences of drone footage where the target is first presented by a telecommunication mast and then by a smoke cloud originating from an industrial chimney. Using these experiments, the study shows that the 3D centre point of the target can be determined both using robust multi-view triangulation or particle filters. The latter of which is also capable of providing a rough estimate of the target object’s shape and the resulting localisation uncertainty.
II Background
The task at hand can be thought of as a form of camera-based multi-view 3D reconstruction. The earlier developments of 3D object reconstruction from multi-view imagery have been well documented by Hartley and Zisserman [Hartley2004]. These methods relied on point correspondence, well-estimated camera parameters, and bundle adjustments. Later developments, such as COLMAP [schoenberger2016sfm] and 3D Gaussian splatting [kerbl3Dgaussians], have greatly improved the structure from motion map generation and 3D modelling, respectively. Unfortunately, the methods still require finding a large number of point correspondences between frames, making them computationally heavy for iterative real-time target localisation using edge devices, which is desirable in, for example, aerial robotic applications.
The 3D reconstruction methods have been poorly studied in the context of smaller and more distant objects. In such situations, any noise in the camera pose estimation causes larger errors in the final reconstruction. This suggests that robust techniques such as Bayesian filters could offer feasible solutions. Besides the camera pose estimation errors, typical modern camera-based sensing solutions leverage neural network models for detecting key features, which complicates keypoint matching between frames in scenarios where we do not want the target object to be reduced to a single point in 3D space.
@@ II Background
In the multisensor context, the use of cameras has been slightly more common as well. However, in the multicamera scenario, the problem is inherently different from that of a single camera, due to the possibility of using the discrepancies between the multiple sensor locations with simultaneous observations. Besides, the metrics in the literature have been focused on the 2D labels as with individual cameras [AMOSA2023126558].
@@ III Materials and Methods
We propose estimating the distant target position using a particle filter. We used simulation and real drone captured footage with GNSS-measured camera positions to evaluate and compare the filter’s performance to that of multi-view triangulation.
@@ III-A Problem Setup
where $\mathbf{y}$ is the projected point in homogeneous coordinates, $\mathbf{K}$ is the intrinsic camera matrix, $\mathbf{M}$ is the extrinsic camera matrix, and $\mathbf{x}$ is the original 3D point in homogeneous coordinates. The homogeneous coordinates use an extra dimension to simplify the matrix operations, such that $\mathbf{y}=(y_{1},y_{2},y_{3})^{\top}$ and $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})^{\top}$, where the pixel coordinates of the projection are:
@@
The false positives in the simulated segments were generated by defining random rectangular sections of the image and setting these pixels equal to the real positive segments. These false-positive rectangles were defined by their height and width in pixel dimensions. They were generated randomly for each image, based on a false positive rate $\rho_{FP}$ and the maximum number of false positives $\mathbf{Max_{FP}}$. Correspondingly, the false positives were removed from the following frames based on a false positive dismissal rate $\delta\rho_{FP}$. Unless the false positive was removed, it was kept for the subsequent images.
@@
III-B Multi-view Triangulation
The simplest solution to the problem is presented by multi-view triangulation. The method allows determining the centre point of the target object by solving a least squares estimate of a point between the camera rays obtained from different viewing points. In practice, this means generating a 3D ray based on each segmentation frame and camera matrix, solving a direct linear transform (DLT), and finding the least squares position.
In practice, we implemented multi-view triangulation for the task by reducing the segments into individual pixel coordinates by taking their centre positions and solving the DLT. The DLT was solved by first generating, for each camera pose and the corresponding segment centre, two linear equations that were used to ensure that the 3D point lies in the correct horizontal and vertical planes correspondingly
$y_{3}(\mathbf{M_{1}\hat{x}})-y_{1}(\mathbf{M_{3}\hat{x}})=0,$ (7)
$y_{2}(\mathbf{M_{3}\hat{x}})-y_{3}(\mathbf{M_{2}\hat{x}})=0,$ (8)
where $\mathbf{M_{1}}$, $\mathbf{M_{2}}$, and $\mathbf{M_{3}}$ denote the rows of the camera matrix $\mathbf{M}$. Stacking all the equations provided a system of linear equations, $\mathbf{A\hat{x}}=0$, which was minimised using the singular value decomposition of $\mathbf{A}$. The minimised solution for $\mathbf{\hat{x}}$ was the 3D target position estimate in homogeneous coordinates.
To make the triangulation more robust to outliers, which are abundant in the studied scenario, we used the random sample consensus (RANSAC) algorithm. The RANSAC discards the outlier camera positions and segments by evaluating a reprojection error
[…]
where $\mathbf{\bar{y}}$ is the reprojection of a candidate position $\mathbf{\hat{x}}$ computed by the standard DLT of two randomly sampled camera positions and segments. $E_{reproj}$ was minimised until a sufficient number ($\geq 80\%$) of observations were considered inliers ($E_{reproj}<2$). If less than 80% were inliers, the configuration with the highest number of inliers was used. Finally, all the inlier positions were used to compute a final estimate using the standard DLT. The algorithm thus discarded outlier segments, finding a more reliable solution.
III-C Particle Filter
The multi-view triangulation was, in this scenario, only able to determine the centre of position of the target object, which is unreliable for a practical task such as wildfire localisation. As such, we also show that the same estimates can be produced using a particle filter, which simultaneously provides rough object shape and uncertainty estimates.
@@ III-C Particle Filter
We defined the filter such that the particles were initialised uniformly along a line in 3D space corresponding to the camera ray of the first observation. The limits of the distribution were set at 50 and 30 000 metres from the camera.
@@
The step was first taken immediately after the initialisation. At each prediction step, the particles were injected with independent Gaussian noise defined by the individual particle variation $\mathbf{\sigma_{p}}$, based on a domain-specific constant. This meant that after each prediction step, the full distribution could be expressed as a new combination of $N$ three-dimensional Gaussians, where $N$ is the number of particles. This distribution serves as the target position prediction of the model.
@@
The update was only done given that positive pixels were observed. In the update step, the pixel distances between the projected particles and the positive observations were compared. Each particle, projected inside the camera frame, was assigned a weight, $\mathbf{\omega}_{p}$, relative to the distance from the positive pixels such that:
$\mathbf{\omega}_{p}=e^{-\min((\mathbf{obs}-\mathbf{p_{proj}})^{2})},$ (10)
where $\mathbf{obs}$ is an array of the positive observations and $\mathbf{p_{proj}}$ is an array where each element presents the same individual projected particle in the pixel coordinates, with the number of elements matching those of the observations.
@@
We used the bootstrap implementation of the particle filter, meaning that the resampling was performed after each update step, by using the weights as the probabilities of drawing the corresponding points from the set of particles. The same number of particles as the previous set was drawn in this manner, resulting in a new set where the highest weighted particles were repeated multiple times. This resulting distribution was then used in the next prediction step, where particles were first updated randomly, effectively doing Gaussian kernel smoothing for the distribution, before projecting them again to the camera frame. We chose the bootstrap version of the particle filter, due to better initial results.
@@ III-D Metrics
The performance of the proposed method was quantified, most importantly, by the root mean square error (RMSE), which was computed between the predicted particles, $\mathbf{p}_{i}$, and the mean of the target location, $\mathbf{m}_{t}$:
$\mathrm{RMSE}=\sqrt{\dfrac{\sum_{i=1}^{n_{p}}(\mathbf{p}_{i}-\mathbf{m}_{t})^{2}}{n_{p}}},$ (11)
where $n_{p}$ is the number of particles. This is equal to the Euclidean distance between the predicted and ground truth means.
Another measure used for evaluating the particle filter, specifically in the simulation, was the ratio of particles which fell in the target object region in the 3D space. The ratio was computed as $ratio=N_{in}/N_{out}$, where $N_{in}$ and $N_{out}$ were the number of particles inside and outside the target object correspondingly. The ratio enabled quantitatively observing how well the particle filter distribution converges to that of the actual target object.
For both metrics, from the simulations, we report the mean between 200 and 1000 metres of camera translation, and for the RMSE, we also report the minimum obtained value during the test. The first 200 metres were dismissed due to being mostly dependent on the initialisation, and 1000 metres was the maximum translation in the experiments. For all results, the metrics were computed over an average of ten simulations to reduce noise caused by the randomness of the experiments. For the particle filter, the multiseed configuration was also used when evaluating with empirical data due to the method’s dependency on the pseudo-randomly varied particles.
@@ III-E Empirical Data
To validate that the performance of the proposed method holds for real-world applications, we tested the system on two drone-captured video sequences. The position of the drone in each was recorded using GNSS, and the approximate geolocations of the targets were known.
To capture the first sequence with a telecom mast target, we used a DJI Matrice 350, equipped with an AR0234 camera and an Applanix APX-15 UAV GNSS. An NVIDIA Jetson Orin NX was used to record the data. We performed no calibration for the camera or the GNSS. This means that for the camera, only the parameters provided by the camera manufacturer were used [arducamAR0234Arducam]. The camera was used to record at full HD (1080 by 1920 pixels) resolution with a horizontal FOV of 90 and vertical FOV of 50.625. This means that the intrinsic parameters were set as $f_{x}=1200$, $f_{y}=1200$, $c_{x}=960$, and $c_{y}=540$. The lens was assumed to cause no distortions. The GNSS antenna was mounted approximately 30 centimetres away from the camera with an external IMU attached to the camera. Neither the boresight nor the lever arm of the mounting was taken into account for the experiments. This means that the camera pose estimation had at least a systematic error in the range of tens of centimetres in addition to any sensor-caused errors. The manufacturer reported RMSE for the position is 0.02 to 0.05 metres, for the roll and pitch 0.025 degrees, and for the heading 0.080 degrees when using differential GNSS [APX15UAV]. The maximum camera translation in the sequence was approximately 250 metres, with the mean distance to the target at approximately 700 metres.
The second sequence, with an industrial smoke cloud as the target, was captured using a DJI Mini 3. Again, the video was captured in full HD resolution but with a slightly different FOV. The DJI Mini 3 horizontal FOV of 82.1° corresponds to a focal length $f_{x}=f_{y}=1102.41$. The reported hovering accuracy of the GNSS was 0.5 metres vertically and 1.5 metres horizontally [DJIMini3]. The angular error ranges of the camera pose estimate were not reported. As for the first sequence, no additional calibration was performed for the system before recording the data. The maximum camera translation in the sequence was approximately 230 metres with the mean distance to the target at 1770 metres.
@@
Method Max $\nu_{rot}$ ($\circ$) Max $\nu_{t}$ (m) $\rho_{FP}$ $\delta\rho_{FP}$ Max FPFP $\rho_{FN}$ $\rho_{PFN}$ $\delta\rho_{PFN}$ RMSE min (m) RMSE 200-1k (m) Ratio 200-1k
MVT 0 0 0 - 0 0 0 - 1.06 3.95 -
MVT 0.5 0.1 0 - 0 0 0 - 1.08 3.79 -
MVT 0.5 0.1 0.1 0.2 3 0 0 - 1.08 147.84 -
MVT 0.5 0.1 0.1 0.2 3 0.1 0 - 1.32 277.88 -
MVT 0.5 0.1 0.1 0.2 3 0.1 0.1 0.2 1.54 257.67 -
RMVT 0 0 0 - 0 0 0 - 1.06 3.92 -
RMVT 0.5 0.1 0 - 0 0 0 - 1.25 3.88 -
RMVT 0.5 0.1 0.1 0.2 3 0 0 - 1.21 3.66 -
RMVT 0.5 0.1 0.1 0.2 3 0.1 0 - 1.26 3.21 -
RMVT 0.5 0.1 0.1 0.2 3 0.1 0.1 0.2 1.20 3.94 -
PF 0 0 0 - 0 0 0 - 9.44 44.83 0.15
PF 0.5 0.1 0 - 0 0 0 - 9.57 45.67 0.15
PF 0.5 0.1 0.1 0.2 3 0 0 - 10.99 46.97 0.15
PF 0.5 0.1 0.1 0.2 3 0.1 0 - 13.04 46.86 0.15
PF 0.5 0.1 0.1 0.2 3 0.1 0.1 0.2 17.87 63.97 0.15
The images of the first sequence were post-processed using a horizontal Sobel operation with image erosion and dilation. This way, we obtained a sequence of binary segments with a known target location and estimated camera positions, corresponding to a drone-based target localisation task. The known target geolocation also enabled measuring the quality of the localisation in ground coordinates.
From the second sequence, the smoke cloud was segmented using SAM 3 [carion2026sam] with the text prompt ’smoke’. The model recognised separate instances of smoke clouds, but only the nearest visible smoke cloud was used in the evaluation. The other segments were discarded. The errors for the test were measured from the top of the industrial chimney. As the smoke cloud was moving away from the chimney, the expected error was non-zero. Based on visual estimation, errors from 50 to 200 metres were expected.